# Fully data-driven inverse hyperelasticity with hyper-network neural ODE fields

### v1

@@ 2.1 Recovering arbitrary heterogeneous material fields from strain field data and boundary conditions
The details of the new solution procedure are illustrated in Fig 1. Note that the goal is to obtain a Piola stress field $\bf{P}(\bf{X})$ that satisfies equilibrium, i.e., $\nabla_{\bf{X}}\cdot{\bf{P}}=\mathbf{0}$ in the domain $\Omega$, along with the prescribed traction boundary conditions. As we explain in the Methodology (Section 5), to find such a stress field, we minimize a loss function containing both terms (referred to as Training loss in Fig 1). The stress field, in turn, is evaluated from ${\bf{P}}=2{\bf{F}}\,\,\partial\Psi(\bf{F})/\partial\bf{C}$, in which the strain density function $\Psi$ is approximated in terms of summation of NODEs (shown as NODE Mat. Model in Fig 1, see also (9)), where the deformation gradient field $\bf{F}$ is either directly provided as input data or calculated from the displacement field data (Fig 1). Because of the divergence operator in the loss function, we need a spatially continuous representation of $\bf{F}$. In this regard, there are two options: interpolating $\bf{u}$ continuously and differentiating to get $\bf{F}$. Alternatively, one can apply a finite difference method, such as the central difference method, on discrete $\bf{u}$, and subsequently interpolate $\bf{F}$ continuously from the discrete field with a neural network. In Fig 1, We assume that displacements are given as the input data. Next, using finite differences we obtain the discretized deformation gradient field. Subsequently, a continuous interpolation of the discrete $\bf{F}$ is obtained from a feed-forward neural network with Fourier features encoding to capture the sharp gradients. The key component of the method is the representation of the strain energy function field $\Psi({\bf{X}},{\bf{F}},\rm\,{cof}{\bf{F}},J)$. We recover a full material model point-wise. In other words, for each $\bf{X}$, we recover a polyconvex constitutive equation $\Psi_{\Theta}(\bf{F},\mathrm{cof}\bf{F},\det\bf{F})$ in terms of summation of NODEs as in our previous work[36]. The notation $\Theta$ is used for the parameters of the NODE model. To discover the heterogeneous material response, the field of parameters $\Theta(\bf{X})$ is obtained from a multi-layer perceptron (MLP), referred to as hyper-network in Fig 1, with learnable parameters $\theta^{\text{hyper}}$. Then, the outcomes are the identified profile of the inhomogeneity and the corresponding stress–strain relations, as shown in Fig. 1 for a square with an X-shaped inhomogeneity. In particular, because the material field is parameterized by $\Theta(\bf{X})$ and it is not a closed-form model, we report the trace of the stress corresponding to a uniform deformation $\lambda_{x}(\mathbf{X})=\lambda_{y}(\mathbf{X})=1.1$, as depicted in Fig. 1 by the blue square with a white X. Also, for two selected points, the stress-stretch response under biaxial tension is shown in Fig 1, showing excellent agreement between ground truth and the NODE model at those locations. To obtain the above results, three sets of function approximators were employed. The interpolation MLP has an architecture of $[81,40,40,4]$, with an initial Fourier feature layer of size $[2,40]$. The NODEs use core MLPs with architecture $[1,4,4,4,1]$. The architecture of the MPL hyper-network is $[80,40,40,126]$, with an initial Fourier feature layer of size $[2,40]$.
@@
We further test out the methodology with a synthetic plate with a $P$-shape inhomogeneity in the middle, as shown in Fig 2. The true material response can be found in Fig 2f. In this case, the inclusion and the surrounding materials are both neo-Hookean with shear moduli $\mu=2,1$ Pa, respectively. The input data consists of deformation gradient field computed using JAX-FEM [38] at selected increments of equi-biaxial loading (see Fig 2a), along with traction boundary conditions. The outcomes of the inverse problem are the identification of the stress field ${\bf{P}}({\bf{X}},t)$ (Fig. 2b) and the characterization of the underlying material response (Figs. 2e and f). While not explicitly shown, the identified stress field ${\bf{P}}({\bf{X}},t)$ in Fig. 2b satisfies the equilibrium condition $\nabla_{\bf X}\cdot{\bf P}({{\mathbf{X}}},t)=0$, which is enforced by minimizing the loss function during the optimization of the hyper-neural network parameters. Satisfaction of boundary load data is shown in Fig 2e. The identified material is not closed form and, as a result, cannot be represented in terms of a heterogeneous modulus field. Rather, Fig 2d shows the trace of the stress for a uniform deformation $\lambda_{x}({\bf X})=\lambda_{y}({\bf X})=1.1$. Note that the stress in Fig 2e is not an equilibrium stress field. Instead, it should be interpreted as an approximation to a modulus field. This is an approximation; in reality, at every point $\bf{X}$, we have a polyconvex strain energy potential, $\Psi({\bf{X}},{\bf{F}},\rm\,{cof}{\bf{F}},J)$. Fig 2f shows the evaluation of the resulting Piola stress from the polyconvex energy for two points, as well as comparison against the true neo-Hookean models. Thus, the proposed algorithm is able to capture an equilibrium stress field and a field of data-driven NODE strain energy potentials, given full-field strains and boundary force data. To generate the above results, an MLP with architecture $[201,40,40,4]$ and an initial Fourier feature layer of size $[2,100]$ was used to interpolate the deformation gradient field. All NODEs share the same architecture as above: $[1,4,4,4,1]$. Additionally, the hyper-network MLP has architecture $[80,40,40,126]$, with an initial Fourier feature layer of size $[2,40]$ as above.
@@ 2.2 Discovering anisotropy
One of the main features of the proposed method is that there is no restriction on the material field except continuity by the nature of $\Theta(\bf{X})$ approximated by the MLP hyper-network. However, there is no specific format existing in closed form, and there is no assumption on the type of material or type of heterogeneity. In this regard, in the synthetic example of Fig 3a we select one region to be isotropic neo-Hookean, and the another region is anisotropic and based on the soft tissue constitutive model proposed by Gasser, Ogden, (GOH) [39]. The sample was subjected to biaxial loading in JAX-FEM, and we report the discretized strain field directly from the finite element simulation in Fig 3c, and the interpolated strains in Fig 3d using the Fourier feature neural network. The loss, as before, consists of linear momentum balance in the static limit and boundary loads. The equilibrium stress field after training is shown in Fig 3e. The trace of the stress at a constant strain field is depicted in Fig 3f to compare against the ground truth material in Fig 3a. Note that the model automatically discovers the anisotropy. Evaluation of the ground truth material behavior against the discovered NODE model at two locations is shown in Fig 3b, confirming that the training naturally leads to anisotropic response prediction in the inclusion, compared to the isotropic response in the surrounding region. In this example, the following architectures were employed: $[201,40,40,4]$ with an initial Fourier feature layer of size $[2,100]$ for interpolating the ground truth deformation gradient field $\mathbf{F}$; $[1,4,4,4,1]$ for the MLP cores used in the NODEs; and $[20,40,40,126]$ with an initial Fourier feature layer of size $[2,10]$ for the MLP hyper-network.
@@ 2.3 Performance in the presence of noise
In the previous synthetic examples, data was assumed to be noise-free. However, we anticipate noise shall have an effect on the final results. Hence, as the next synthetic example, we illustrate this fact by introducing noise, which can be interpreted as the error of measurement in experimental data. To this end, we consider the case where the noise is introduced in displacement measurement, i.e., ${{\bf{u}}^{{\rm{noise}}}}={\bf{u}}+{\bf{\epsilon}}$ with ${\bf{\epsilon}}$ denoting Gaussian noise with mean and variance $(\mu,\sigma)$. In this example, we consistently assume $\mu=0$ and three cases of noise: noiseless, $\sigma=0.0007$, and $\sigma=0.001$. Moreover, we consider a square plate consisting of a 50by $5050\,\rm{by}\,$ 50 quadrilateral mesh with a ring-type feature. Additionally, we consider five steps of uniform loading with $\lambda_{steps}=[1.05544,1.07795,1.10065,1.12326,1.14587]$. Fig 4a shows the generating material and corresponding displacement fields from the finite element simulation at $\lambda({\bf X})=1.14587$. Fig 4b show the strain field interpolated by a neural network with Fourier features after injecting different levels of noise to the displacement field. The corresponding error with respect to ground truth data increases by increasing $\sigma$, as expected, but with errors in the strain being an order of magnitude larger than those in the displacements. Fig 4d shows the result of the inverse problem. The inverse analysis is sensitive to the noise in the data. In agreement with previous examples, the inverse problem works well in the noiseless case, but the method failed to accurately recognize the non-homogeneity for both $\sigma=0.001,0.0007$. We remark that the noise of $\sigma=0.0007$ means that the 99 percent confidence interval of the displacement variation with respect to ground truth $[-0.001806,0.001806]$, while the noiseless displacements are on order 0.150.15~{}0.150.15. Note also that this error leads to $6$% error in $\bf{F}$, as shown in Fig. 4c. Thus, about $0.06$ error is tolerated in the non-dimensional input $\bf{F}$. This is the metric that should be assessed to anticipate whether or not the inverse problem is likely to work accurately.
@@ 5.1 Inverse problem setup
It can be shown that any anisotropic strain density function defined above is characterized by an isotropic function of the right Cauchy-Green strain and the corresponding structural tensors ([56, 55]): ${{\hat{\Psi}}^{{\rm{aniso}}}}({\bf{C}},{\bf{X}})={{\hat{\Psi}}^{{\rm{iso}}}}({\bf{C}},{{\bf{L}}_{1}},...,{{\bf{L}}_{n}},{\bf{X}})$. From this point onward, for the sake of brevity, we denote isotropic functions with $\Psi$. The next concept, which has a mathematical background rather than a physical one, is the notion of polyconvexity [57]. Specifically, this property ensures that the strain energy satisfies the ellipticity of the acoustic tensor [58], and thus, the displacement wave speed is always a real number— which, together with the growth condition, ensures the existence of a global minimizer. Mathematically, a strain density energy is polyconvex if and only if $\Psi({\bf{F}},{\bf{X}})={\Psi^{\mathcal{P}}}({\bf{F}},\,{\rm{cof}}\,{\bf{F}},\,\det\,{\bf{F}},{\bf{X}})$, where $\Psi^{\mathcal{P}}$ is convex with respect to first three arguments [57, 59]. Finally, in the language of data-driven methods, a constitutive is called thermodynamically consistent if (2) holds. Also, it is worth noting that (1)2italic-(1subscriptitalic-)$_{2}$ is satisfied if we consider the following:
@@ 5.4 Synthetic data generation
To generate synthetic data, we used JAX-FEM [38]. We considered the linear shape functions, and the JAX-FEM built-in function was used to calculate deformation gradients. Also, the geometry is a mesh of 50by  $5050\,\,\rm{by}\,\,$ 50 quadrilateral elements for a square with a dimension of unity.
@@ 5.5 Experimental data generation
To generate experimental data, a heterogeneous sample was 3D-printed using a Stratasys PolyJet J750 Digital Anatomy printer (Stratasys, Eden Prairie, MN, USA). The geometry of the sample was taken directly from the MNIST dataset, and sample dimensions were $40\times 40\times 2$ mm. The sample consisted of two components: an inclusion and the surrounding material. The inclusion was composed of FLX9870 with a tangent modulus of about 1400 kPa, while the surrounding material was composed of Agilus30 with a tangent modulus of about 550 kPa. Additionally, to ensure a secure attachment between the sample and the mechanical testing device, rigid clamps were incorporated into the sample design.
Following printing, the sample was prepared for DIC by applying a speckle pattern to one side. A thin base coat of white acrylic ink was first applied, followed by a distribution of small black dots to create a high-contrast pattern. The sample was then mounted onto a uniaxial tensile tester (Instron, Norwood, MA, USA) using pneumatic grips and extended until failure at a rate of 0.2 mm/s. A 100 N load cell (accuracy of $\pm$ 0.2 N) was used to measure force, while displacement and time data were recorded via Instron software. Furthermore, a custom LabVIEW program (National Instruments, Austin, TX, USA, Version 2021 SP1) was used to take images of the sample during extension at a rate of 5 Hz. These images were exported to LaVision DaVis (LaVision, Göttingen, Germany, Version 11) for post-processing. Using the strain tracking feature, displacement fields were captured for a $42\times 41$ grid covering the entire sample area.

### v2

@@ 2.1 Recovering arbitrary heterogeneous material fields from strain field data and boundary conditions
The details of the new solution procedure are illustrated in Fig 1. Note that the goal is to obtain a Piola stress field $\bf{P}(\bf{X})$ that satisfies equilibrium, i.e., $\nabla_{\bf{X}}\cdot{\bf{P}}=\mathbf{0}$ in the domain $\Omega$, along with the prescribed traction boundary conditions. As we explain in the Methodology (Section 5), to find such a stress field, we minimize a loss function containing both terms (referred to as Training loss in Fig 1). The stress field, in turn, is evaluated from ${\bf{P}}=2{\bf{F}}\,\,\partial\Psi(\bf{F})/\partial\bf{C}$, in which the strain density function $\Psi$ is approximated in terms of summation of NODEs (shown as NODE Mat. Model in Fig 1, see also (10)), where the deformation gradient field $\bf{F}$ is either directly provided as input data or calculated from the displacement field data (Fig 1). Because of the divergence operator in the loss function, we need a spatially continuous representation of $\bf{F}$. In this regard, there are two options: interpolating $\bf{u}$ continuously and differentiating to get $\bf{F}$. Alternatively, one can apply a finite difference method, such as the central difference method, on discrete $\bf{u}$, and subsequently interpolate $\bf{F}$ continuously from the discrete field with a neural network. In Fig 1, We assume that displacements are given as the input data. Next, using finite differences we obtain the discretized deformation gradient field. Subsequently, a continuous interpolation of the discrete $\bf{F}$ is obtained from a feed-forward neural network with Fourier features encoding to capture the sharp gradients. The key component of the method is the representation of the strain energy function field $\Psi({\bf{X}},{\bf{F}},\rm\,{cof}{\bf{F}},J)$. We recover a full material model point-wise. In other words, for each $\bf{X}$, we recover a polyconvex constitutive equation $\Psi_{\Theta}(\bf{F},\mathrm{cof}\bf{F},\det\bf{F})$ in terms of summation of NODEs as in our previous work[36]. The notation $\Theta$ is used for the parameters of the NODE model. To discover the heterogeneous material response, the field of parameters $\Theta(\bf{X})$ is obtained from a multi-layer perceptron (MLP), referred to as hyper-network in Fig 1, with learnable parameters $\theta^{\text{hyper}}$. Then, the outcomes are the identified profile of the inhomogeneity and the corresponding stress–strain relations, as shown in Fig 1 for a square with an X-shaped inhomogeneity. In particular, because the material field is parameterized by $\Theta(\bf{X})$ and it is not a closed-form model, we report the trace of the stress corresponding to a uniform deformation $\lambda_{x}(\mathbf{X})=\lambda_{y}(\mathbf{X})=1.1$, as depicted in Fig 1 by the blue square with a white X. Also, for two selected points, the stress-stretch response under biaxial tension is shown in Fig 1, showing excellent agreement between ground truth and the NODE model at those locations. To obtain the above results, three sets of function approximators were employed. The interpolation MLP has an architecture of [81,40,40,4][81,40,40,4][ 81 , 40 , 40 , 4 ], with an initial Fourier feature layer of size [2,40][2,40][ 2 , 40 ]. The NODEs use core MLPs with architecture [1,4,4,4,1][1,4,4,4,1][ 1 , 4 , 4 , 4 , 1 ]. The architecture of the MPL hyper-network is [80,40,40,126][80,40,40,126][ 80 , 40 , 40 , 126 ], with an initial Fourier feature layer of size [2,40][2,40][ 2 , 40 ].
@@
We further test out the methodology with a synthetic plate with a $P$-shape inhomogeneity in the middle, as shown in Fig 2. The true material response can be found in Fig 2f. In this case, the inclusion and the surrounding materials are both neo-Hookean with shear moduli $\mu=2,1$ Pa, respectively. The input data consists of deformation gradient field computed using JAX-FEM [38] at selected increments of equi-biaxial loading (see Fig 2a), along with traction boundary conditions. The outcomes of the inverse problem are the identification of the stress field ${\bf{P}}({\bf{X}},t)$ (Fig 2b) and the characterization of the underlying material response (Figs. 2e and f). While not explicitly shown, the identified stress field ${\bf{P}}({\bf{X}},t)$ in Fig 2b satisfies the equilibrium condition $\nabla_{\bf X}\cdot{\bf P}({{\mathbf{X}}},t)=0$, which is enforced by minimizing the loss function during the optimization of the hyper-neural network parameters. Satisfaction of boundary load data is shown in Fig 2e. The identified material is not closed form and, as a result, cannot be represented in terms of a heterogeneous modulus field. Rather, Fig 2d shows the trace of the stress for a uniform deformation $\lambda_{x}({\bf X})=\lambda_{y}({\bf X})=1.1$. Note that the stress in Fig 2e is not an equilibrium stress field. Instead, it should be interpreted as an approximation to a modulus field. This is an approximation; in reality, at every point $\bf{X}$, we have a polyconvex strain energy potential, $\Psi({\bf{X}},{\bf{F}},\rm\,{cof}{\bf{F}},J)$. Fig 2f shows the evaluation of the resulting Piola stress from the polyconvex energy for two points, as well as comparison against the true neo-Hookean models. Thus, the proposed algorithm is able to capture an equilibrium stress field and a field of data-driven NODE strain energy potentials, given full-field strains and boundary force data. To generate the above results, an MLP with architecture [201,40,40,4][201,40,40,4][ 201 , 40 , 40 , 4 ] and an initial Fourier feature layer of size [2,100][2,100][ 2 , 100 ] was used to interpolate the deformation gradient field. All NODEs share the same architecture as above: [1,4,4,4,1][1,4,4,4,1][ 1 , 4 , 4 , 4 , 1 ]. Additionally, the hyper-network MLP has architecture [80,40,40,126][80,40,40,126][ 80 , 40 , 40 , 126 ], with an initial Fourier feature layer of size [2,40][2,40][ 2 , 40 ] as above.
@@ 2.2 Discovering anisotropy
One of the main features of the proposed method is that there is no restriction on the material field except continuity by the nature of $\Theta(\bf{X})$ approximated by the MLP hyper-network. However, there is no specific format existing in closed form, and there is no assumption on the type of material or type of heterogeneity. In this regard, in the synthetic example of Fig 3a we select one region to be isotropic neo-Hookean, and the another region is anisotropic and based on the soft tissue constitutive model proposed by Gasser, Ogden, (GOH) [39]. The sample was subjected to biaxial loading in JAX-FEM, and we report the discretized strain field directly from the finite element simulation in Fig 3c, and the interpolated strains in Fig 3d using the Fourier feature neural network. The loss, as before, consists of linear momentum balance in the static limit and boundary loads. The equilibrium stress field after training is shown in Fig 3e. The trace of the stress at a constant strain field is depicted in Fig 3f to compare against the ground truth material in Fig 3a. Note that the model automatically discovers the anisotropy. Evaluation of the ground truth material behavior against the discovered NODE model at two locations is shown in Fig 3b, confirming that the training naturally leads to anisotropic response prediction in the inclusion, compared to the isotropic response in the surrounding region. In this example, the following architectures were employed: [201,40,40,4][201,40,40,4][ 201 , 40 , 40 , 4 ] with an initial Fourier feature layer of size [2,100][2,100][ 2 , 100 ] for interpolating the ground truth deformation gradient field $\mathbf{F}$; [1,4,4,4,1][1,4,4,4,1][ 1 , 4 , 4 , 4 , 1 ] for the MLP cores used in the NODEs; and [20,40,40,126][20,40,40,126][ 20 , 40 , 40 , 126 ] with an initial Fourier feature layer of size [2,10][2,10][ 2 , 10 ] for the MLP hyper-network.
@@ 2.3 Performance in the presence of noise
[…]
The computed SNR values corresponding to the results in Fig 4b are 19.546619.546619.5466 and 18.943418.943418.9434 for the cases with $\sigma=0.0007$ and $\sigma=0.001$, respectively.
@@ 5.1 Inverse problem setup
It can be shown that any anisotropic strain density function defined above is characterized by an isotropic function of the right Cauchy-Green strain and the corresponding structural tensors ([56, 55]): ${{\hat{\Psi}}^{{\rm{aniso}}}}({\bf{C}},{\bf{X}})={{\hat{\Psi}}^{{\rm{iso}}}}({\bf{C}},{{\bf{L}}_{1}},...,{{\bf{L}}_{n}},{\bf{X}})$. From this point onward, for the sake of brevity, we denote isotropic functions with $\Psi$. The next concept, which has a mathematical background rather than a physical one, is the notion of polyconvexity [57]. Specifically, this property ensures that the strain energy satisfies the ellipticity of the acoustic tensor [58], and thus, the displacement wave speed is always a real number— which, together with the growth condition, ensures the existence of a global minimizer. Mathematically, a strain density energy is polyconvex if and only if $\Psi({\bf{F}},{\bf{X}})={\Psi^{\mathcal{P}}}({\bf{F}},\,{\rm{cof}}\,{\bf{F}},\,\det\,{\bf{F}},{\bf{X}})$, where $\Psi^{\mathcal{P}}$ is convex with respect to first three arguments [57, 59]. Finally, in the language of data-driven methods, a constitutive is called thermodynamically consistent if (3) holds. Also, it is worth noting that ($_{2}$ is satisfied if we consider the following:
@@ 5.4 Synthetic data generation
[…]
@@ 5.5 Experimental data generation
To generate experimental data, a heterogeneous sample was 3D-printed using a Stratasys PolyJet J750 Digital Anatomy printer (Stratasys, Eden Prairie, MN, USA). The geometry of the sample was taken directly from the MNIST dataset, and sample dimensions were 40×40×240$\times$ 40$\times$ 240 × 40 × 2 mm. The sample consisted of two components: an inclusion and the surrounding material. The inclusion was composed of FLX9870 with a tangent modulus of about 1400 kPa, while the surrounding material was composed of Agilus30 with a tangent modulus of about 550 kPa. Additionally, to ensure a secure attachment between the sample and the mechanical testing device, rigid clamps were incorporated into the sample design.
Following printing, the sample was prepared for DIC by applying a speckle pattern to one side. A thin base coat of white acrylic ink was first applied, followed by a distribution of small black dots to create a high-contrast pattern. The sample was then mounted onto a uniaxial tensile tester (Instron, Norwood, MA, USA) using pneumatic grips and extended until failure at a rate of 0.2 mm/s. A 100 N load cell (accuracy of $\pm$ 0.2 N) was used to measure force, while displacement and time data were recorded via Instron software. Furthermore, a custom LabVIEW program (National Instruments, Austin, TX, USA, Version 2021 SP1) was used to take images of the sample during extension at a rate of 5 Hz. These images were exported to LaVision DaVis (LaVision, Göttingen, Germany, Version 11) for post-processing. Using the strain tracking feature, displacement fields were captured for a 42×4142$\times$ 4142 × 41 grid covering the entire sample area.
6 Acknowledgements
ABT acknowledges support of ARO under award W911NF-24-1-0244.
@@ 8 Supplementary material
9 Appendix
9.1 Ablation study
In this section, we present an ablation study of the proposed method. In this regard, we consider the example with an embedded P-shape inhomogeneity. As discussed in Section5, the main components of the method include interpolation of the ground truth deformation gradient, a spatially dependent central NODE, and a hyper-network that defines the spatial variation of the NODE parameters across the domain. Hence, the ablation study we consider in this work is to investigate the effect of the above-mentioned components. For clarity, the ground truth data is shown in the first panel of each figure. To emphasize the differences, we consistently assess the accuracy of each configuration in identifying the inhomogeneity profile embedded in the square plate. To this end, we examine the nominal stress component ${{\bf{P}}}_{xX}$ evaluated under four levels of imposed strain: $\lambda=[1.05,1.10,1.15,1.20]$. Moreover, the central architectures used in this ablation study are as follows: for interpolation of the deformation gradient, the central architecture is [201, 40, 40, 4] with Fourier features [2, 100]; for the NODEs, the central architecture is [1, 4, 4, 4, 1]; and for the hyper-network, the central architecture is [80, 40, 40, 126] with Fourier features [2, 40].
To obtain ${\bf{F}}^{\rm{int}}$, the key factor is incorporating Fourier features, which are essential for capturing sharp variations in the field. Accordingly, we present results for three architectures: [2,10] (Fig 6a), [2,40] (Fig 6b), and [2,100] (Fig 6c). These figures clearly highlight the effect of Fourier features. In particular, increasing the Fourier feature parameter enhances the resolution of the embedded inhomogeneity profile.
The next component of the proposed method we consider is the architecture of the NODE. In this regard, we define three cases: a coarse architecture [1, 2, 1], a finer one [1, 4, 4, 4, 1], and the finest [1, 8, 8, 8, 1]. The corresponding results are shown in Fig 7b, Fig 7c, and Fig 7d, respectively, with the ground truth data shown in Fig 7a. As can be seen, the architecture [1, 4, 4, 4, 1] provides a better identification of the inhomogeneity profile, slightly outperforming [1, 8, 8, 8, 1]. This suggests that over-parameterization of the NODE may have an adverse effect on the performance of the inverse problem.
Finally, we investigate the effect of different hyper-network architectures on material identification using the proposed method. To this end, we consider three configurations: a coarse architecture [20, 10, 126] with Fourier features [2, 10], a finer architecture [80, 40, 40, 126] with Fourier features [2, 40], and the finest [80, 60, 60, 126] with Fourier features [2, 80]. The corresponding results are shown in Fig 8b, Fig 8c, and Fig 8d, respectively. Similar to the NODE architecture, an over-parameterization effect is observed in the identification of the inhomogeneity profile when using the most complex hyper-network architecture. This observation suggests that beyond a certain complexity, additional network depth may not yield significant improvements in inhomogeneity profile identification.
9.2 Method Breakdown
We present detailed results of the proposed method for the example with an embedded X-shaped inhomogeneity, as illustrated in Fig 9. Fig 9a shows the outcome of the first training phase, where a homogenized field is learned. Fig 9b displays the homogenized field after training the hyper-network to reproduce the NODE parameters. As shown, the hyper-network successfully captures the NODE parameters and accurately reconstructs the homogenized field. Subsequently, Fig 9c presents the results of the second pre-training step, in which the hyper-network is trained to define a NODE field that yields average stresses corresponding to a non-homogeneous strain field. Finally, the stress field obtained from the final training phase is shown in Fig.9d, alongside the ground truth equilibrium stress field in Fig. 9e for validation. As observed, the proposed method yields results in good agreement with the ground truth.
Figure 6: Ablation study to demonstrate the effect of Fourier features used in the interpolation of ground truth data on the performance of the new inverse method. The ground truth stresses under constant strain fields are shown in (a). The corresponding results predicted by the method, using Fourier features with architectures [2, 10], [2, 40], and [2, 100] for interpolation, are shown in (b), (c), and (d), respectively.
Figure 7: Ablation study to show the effect of the architecture used in the core NODEs. The ground truth stresses under constant strain fields are shown in (a). Similar results corresponding to the architectures [1, 2, 1], [1, 4, 4, 4, 1], and [1, 8, 8, 8, 1] are shown in (b), (c), and (d), respectively.
Figure 8: Ablation study to indicate how various architecture used in hyper-network can affect the final outcome of the inverse problem. The ground truth stresses under constant strain fields are shown in (a). Panels (b), (c), and (d) show the corresponding results obtained using the architectures [20, 10, 126] with Fourier features [2, 10], [80, 40, 40, 126] with Fourier features [2, 40], and [80, 60, 60, 126] with Fourier features [2, 40], respectively.
Figure 9: An illustrative example showing the results of the major steps in the proposed inverse method. Panel (a) presents the results after training a homogeneous NODE to recover the average stress field. Panel (b) shows the results after training the hyper-network to generate the parameters of the homogeneous NODE. The stress field resulting from a non-homogeneous deformation gradient after the second pre-training is shown in panel (c). Panel (d) presents the final result after the second pre-training. The ground truth field corresponding to panel (d) is also shown for comparison in panel (e).